OpenStudy (anonymous):

What is the minimum and maximum value of y=-3+4 cos[5pi/6(x+4)]? @mathmale @ranga

zepdrix (zepdrix):

\[\Large\bf\sf y=-3+4\cos\left[\frac{5\pi}{6}(x+4)\right]\]So we need critical points and stuff. Have you tried taking a derivative yet?

OpenStudy (anonymous):

I have not learned that yet

zepdrix (zepdrix):

Oh I'm getting ahead of myself :) sorry bout that hehe

OpenStudy (anonymous):

nay the largest value of cosine is 1 the largest value of 4 times cosine is 4 the largest value of -3 plus 4 cosine is 4 - 3 = 1